OpenStudy (anonymous):

could you help me with these question @mayankdevnani 47. Solve 5x = 15 for x. (Points : 3) 0 1 2 3 48. Solve x + 12 = 3 for x. (Points : 3) -15 -9 9 15 49. Solve 3x + 7 = –5 for x. (Points : 3) 3 4 -3 -4

OpenStudy (mayankdevnani):

47)\[\huge 5 \times x=15\] then, divide 5 on both the sides, \[\huge \frac{5 \times x}{5}=\frac{15}{5}\] so what we get??? @jayfer

OpenStudy (mayankdevnani):

@jayfer reply....

OpenStudy (anonymous):

could you just give me the awnsers cuz its late and i really need this study guide

OpenStudy (mayankdevnani):

ok....

OpenStudy (mayankdevnani):

can i give direct answers?

OpenStudy (anonymous):

ya.

OpenStudy (mayankdevnani):

47) 3 48) -9 49) -4

OpenStudy (mayankdevnani):

ok... @jayfer i completed your wish!!!

OpenStudy (anonymous):

could you help me with a few more @mayankdevnani